# Artificial Intelligence Exploring the Patent Field

### v1

@@ 6 Evaluation Metrics
Evaluation metrics quantify a model’s performance, offering a clear and objective measure of how well the model can achieve the intended task. This is essential for understanding the model’s capabilities and limitations. In addition, metrics allow researchers and developers to compare the effectiveness of different models or algorithms on the same task. By understanding the most relevant metrics to a specific task, developers can select or design models that are optimized for those metrics. Different metrics are used based on task types. We introduce some common evaluation metrics for different tasks, including classification (Section 6.1), information retrieval (Section 6.2), and text generation (Section 6.3).
6.1 Classification
@@ 6.2 Information Retrieval
usually serve for evaluation [103].
$N_{R}$ is the number of relevant documents retrieved, $N_{T}$ the total number of documents retrieved, and $N_{K}$ the total number of real relevant documents. $Q$ denotes the total number of required queries, $m_{q}$ the number of relevant documents for the q-th query, $n_{q}$ the number of documents retrieved for the q-th query, $P(k)$ the precision at the k-th position, and $\sigma(k)$ an indicator function that equals 1 if the k-th document is relevant, and 0 otherwise.
@@
is a metric specifically designed to evaluate the performance of patent retrieval systems [104]. $N$ represents the number of relevant documents, $N_{max}$ the maximum number of documents returned by the retrieval system, and $r_{i}$ the rank of the i-th relevant document in the retrieval results. It calculates the average deviation of the actual ranking of relevant documents from their ideal ranking with normalization.
6.3 Text Generation
@@
Researchers have developed automated evaluation methods for text generation to improve efficiency and reduce costs. These approaches compare generated texts with referenced texts (gold standard) to obtain the performance score. We briefly introduce four of the most commonly used methods, including BLEU [105], ROUGE [106], METEOR [107], and BERTScore [108].
@@
Bilingual evaluation understudy (BLEU) score [105] quantifies how much the generated texts match the high-quality reference texts, by comparing the n-grams to count the number of exact matches. BLEU is simple and easy to use, but it only focuses on literal matches, overlooking overall sentence semantics. BLEU was originally designed for machine translation but could also be used for other generation tasks.
@@
Recall-oriented understudy for gisting evaluation (ROUGE) counts the number of overlapping units, such as n-gram and word sequences, to obtain the performance score [106]. While BLEU focuses more on precision, i.e., how many generated texts match reference texts, ROUGE prioritizes more recall performance, i.e., how many reference texts are covered by generated texts. Hence, this method is more suitable for text summarization tasks. Similar to BLEU, it evaluates text content, without considering semantics or synonyms.
@@
Metric for evaluation of translation with explicit ordering (METEOR) improves on BLEU by accounting for both exact word matches and similar words based on stemming and synonyms [107]. Thus, METEOR can better account for semantic variations compared to BLEU. In addition, METEOR provides a more balanced assessment by considering both precision and recall. However, METEOR is more complex and needs adjustments for different tasks.
@@
BERTScore [108] leverages the contextual embedding from BERT [43] and calculates cosine similarity between output text embedding and referenced embedding. It can effectively capture complex and subtle semantic information but requires significant computational resources for evaluation.

### v2

@@ Appendix B Evaluation Metrics
Evaluation metrics quantify a model’s performance, offering a clear and objective measure of how well the model can achieve the intended task. This is essential for understanding the model’s capabilities and limitations. In addition, metrics allow researchers and developers to compare the effectiveness of different models or algorithms on the same task. By understanding the most relevant metrics to a specific task, developers can select or design models that are optimized for those metrics. Different metrics are used based on task types. We introduce some common evaluation metrics for different tasks, including classification (Section B.1), information retrieval (Section B.2), and text generation (Section B.3).
B.1 Classification
@@ B.2 Information Retrieval
usually serve for evaluation (Baeza-Yates et al., 1999).
$N_{R}$ is the number of relevant documents retrieved, $N_{T}$ the total number of documents retrieved, and $N_{K}$ the total number of really relevant documents. $Q$ denotes the total number of required queries, $m_{q}$ the number of relevant documents for the q-th query, $n_{q}$ the number of documents retrieved for the q-th query, $P(k)$ the precision at the k-th position, and $\sigma(k)$ an indicator function that equals 1 if the k-th document is relevant, and 0 otherwise.
@@
is a metric specifically designed to evaluate the performance of patent retrieval systems (Magdy and Jones, 2010). $N$ represents the number of relevant documents, $N_{max}$ the maximum number of documents returned by the retrieval system, and $r_{i}$ the rank of the i-th relevant document in the retrieval results. It calculates the average deviation of the actual ranking of relevant documents from their ideal ranking with normalization.
B.3 Text Generation
@@
Researchers have developed automated evaluation methods for text generation to improve efficiency and reduce costs. These approaches compare generated texts with referenced texts (gold standard) to obtain the performance score. We briefly introduce four of the most commonly used methods, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2019).
@@
Bilingual evaluation understudy (BLEU) score (Papineni et al., 2002) quantifies how much the generated texts match the high-quality reference texts, by comparing the n-grams to count the number of exact matches. BLEU is simple and easy to use, but it only focuses on literal matches, overlooking overall sentence semantics. BLEU was originally designed for machine translation but could also be used for other generation tasks.
@@
Recall-oriented understudy for gisting evaluation (ROUGE) counts the number of overlapping units, such as n-gram and word sequences, to obtain the performance score (Lin, 2004). While BLEU focuses more on precision, i.e., how many generated texts match reference texts, ROUGE prioritizes more recall performance, i.e., how many reference texts are covered by generated texts. Hence, this method is more suitable for text summarization tasks. Similar to BLEU, it evaluates text content, without considering semantics or synonyms.
@@
Metric for evaluation of translation with explicit ordering (METEOR) improves on BLEU by accounting for both exact word matches and similar words based on stemming and synonyms (Banerjee and Lavie, 2005). Thus, METEOR can better account for semantic variations compared to BLEU. In addition, METEOR provides a more balanced assessment by considering both precision and recall. However, METEOR is more complex and needs adjustments for different tasks.
@@
BERTScore (Zhang et al., 2019) leverages the contextual embedding from BERT (Devlin et al., 2018) and calculates cosine similarity between output text embedding and referenced embedding. It can effectively capture complex and subtle semantic information but requires significant computational resources for evaluation.